# Test-Time 3D Occupancy Prediction

### v3

@@ 5 Conclusion
[…]
@@ A.1 Depth Estimation and Triangulation-Based Calibration for TT-OccCamera
When the outputs of 3DVFMs are already in a metric scale (such as MapAnything [keetha2025mapanything], they can be directly applied. However, for models that do not provide metric outputs (such as VGGT [vggt]), additional calibration is required to align them with the ground-truth occupancy. Using VGGT as an example, we describe below the depth estimation process and a triangulation-based calibration method. VGGT is a feed-forward neural network capable of predicting depth maps and tracking 2D keypoints across frames. We input six surrounding camera views into VGGT to generate per-view depth predictions. Following the original VGGT setup, the input images are resized to a resolution of 294 $\times$ 518. Although VGGT produces consistent and high-quality depth estimates across views, the predictions are in an unscaled unit space and do not correspond directly to real-world metric distances. To address this limitation, we leverage VGGT’s built-in 2D point tracking functionality across multiple views at the same time step. Specifically, we select three adjacent cameras including front, front-left, and front-right, and use VGGT to track sparse 2D keypoints across them. By filtering out low-quality matches using the predicted visibility and confidence scores, we obtain reliable point correspondences between camera pairs, as illustrated in Fig. 7. We then triangulate these matched 2D points using the ground-truth camera intrinsics and extrinsics provided by the dataset, resulting in a sparse but metrically accurate 3D point cloud. Finally, we compare the magnitudes of the triangulated 3D points with those reconstructed from the predicted depth maps at the corresponding image locations, and compute a global scaling factor to align the depth predictions with real-world scale. An example of the final scaled depth prediction is shown in Fig. 8.
@@ A.2 Open-Vocabulary Semantic Segmentation
We now describe the process of open-vocabulary semantic segmentation using OpenSeeD [OpenSeeD] as an example. We adopt OpenSeeD primarily to ensure a fair comparison with SelfOcc [SelfOcc]. Nevertheless, our system is loosely coupled with VFMs and fully compatible with more advanced ones, such as REX-Omni [rexomni]. As shown in Fig. 9, OpenSeeD’s predictions often exhibit noisy and unclear boundaries. Since our focus is not prompt engineering and to ensure a fair comparison with SelfOcc [SelfOcc], we adopt the same query set including: "barrier", "bicycle", "bus", "car", "construction_vehicle", "crane", "motorcycle", "person", "traffic_cone", "trailer", "trailer_truck", "truck", "road", "sidewalk", "terrain", "grass", "building", "wall", "tree", "sky".
@@ A.3 Tracking with RAFT for TT-OccCamera
For $\operatorname{TT\text{-}OccCamera}$, we estimate the optical flow $F_{\text{opt}}$ between two consecutive frames from the same camera using RAFT [raft]. We then compute the ego-motion-induced flow $F_{\text{ego}}$ based on the ground-truth camera intrinsics and extrinsics of the adjacent frames, along with the predicted depth from 3DVFMs. By subtracting the ego flow from the observed optical flow, we obtain the dynamic flow $F_{\text{dyn}}=F_{\text{opt}}-F_{\text{ego}}$, which theoretically captures the motion of dynamic objects in the environment. Although this 2D dynamic flow could, in principle, guide the 3D motion of dynamic Gaussians, back-projecting it into 3D space tends to amplify errors from RAFT and 3DVFMs, resulting in unstable Gaussian motion. To mitigate this, we adopt a compromise strategy by thresholding the dynamic flow magnitude to obtain a dynamic mask that identifies likely moving regions. In the ideal case, a simple thresholding on the magnitude of $F_{\text{dyn}}$ would yield a reliable binary mask for dynamic regions. However, since both $F_{\text{opt}}$ and $F_{\text{ego}}$ are derived from 2D estimations and are subject to noise and inaccuracies, the resulting $F_{\text{dyn}}$ is often highly unreliable and noisy. To further refine the dynamic flow, we leverage the the cues from semantic segmentation models, which provides relatively cleaner object boundaries, to refine the dynamic flow magnitude map. As illustrated in Fig. 10, the raw dynamic flow is noisy, and thresholding it directly often produces fragmented masks that do not correspond to coherent objects. After incorporating instance masks from segmentation models, high-magnitude errors on the background are suppressed, and the resulting dynamic masks become more object-aligned, either an entire object is identified as dynamic or it is not, effectively eliminating partial or spurious activations. The corresponding 3D Gaussians projected onto these regions are treated as dynamic and excluded from static accumulation in the next frame. While this approach does not allow accumulation of dynamic objects as in the LiDAR-based variant, it effectively reduces artifacts caused by noisy motion cues and temporal inconsistencies.
@@ A.4 Tracking with LiDAR for TT-OccLiDAR
Tracking in $\operatorname{TT\text{-}OccLiDAR}$ is generally more reliable than in $\operatorname{TT\text{-}OccCamera}$, as LiDAR point clouds provide more accurate and consistent geometric information. We follow a straightforward strategy: cluster first, then align via ICP. First, we project LiDAR points onto the instance masks predicted by the segmentation model, thereby associating each point with a specific foreground object. Due to the often imprecise boundaries of predicted masks, the resulting instance-level point sets can contain substantial noise. To address this, we apply DBSCAN clustering [dbscan] to each instance’s point cloud to extract its core structure and eliminate outliers. This approach proves effective in significantly removing noise, as illustrated in the left column of Fig. 11, where gray points are obtained by directing projecting onto segmentation masks and green points represent the denoised output after DBSCAN clustering (slightly translated for observation). We then perform object-level matching across adjacent frames based on the spatial proximity and shape similarity of the filtered point clusters. For each matched pair, the 3D flow is estimated using the Iterative Closest Point (ICP) algorithm [ICP]. Qualitative results are presented in the right column of Fig. 11, where green, blue, and red points represent the source points, destination points, and the ICP-transformed source points, respectively. Green arrows indicate the estimated 3D flow vectors. The effectiveness of the ICP-based alignment can be clearly observed. Finally, matched points are propagated to the next frame, while unmatched instances from the previous frame are discarded to avoid the accumulation of errors caused by moving or disappearing objects.
@@ B.1 Performance under Challenging Conditions
A key strength of our system is that it is built upon large-scale foundation models, which are trained on diverse datasets and therefore exhibit strong generalization beyond specific domains. Leveraging these models gives TT-Occ a natural advantage over task-specific approaches such as SelfOcc [SelfOcc], particularly under challenging conditions. To assess robustness, we evaluated both the camera- and LiDAR-based variants on nighttime and rainy scenes from the nuScenes [nuScenes] dataset. As shown in Table 6, TT-Occ consistently surpasses SelfOcc across all challenging scenarios, demonstrating improved accuracy and robustness.
Table 6: Robustness evaluation on rainy and nighttime nuScenes [nuScenes] scenes (mIoU). R.: rainy. N.: nighttime. TT-Occ variants consistently outperform the task-specific SelfOcc [SelfOcc] model across all challenging conditions.
@@
T-Camera 21.2 13.6 11.2 9.4 10.3 13.1
T-LiDAR 27.0 18.9 12.0 13.9 16.2 17.6

### v4

@@ 5 Conclusion
6 Acknowledgment
This work was partially supported by ARC DE240100105, DP240101814, DP230101196, BA24006, and ARC Industrial Transformation Research Hubs IH230100013.
@@ A.1 Depth Estimation and Triangulation-Based Calibration for TT-OccCamera
When the outputs of 3DVFMs are already in a metric scale (such as MapAnything [19], they can be directly applied. However, for models that do not provide metric outputs (such as VGGT [37]), additional calibration is required to align them with the ground-truth occupancy. Using VGGT as an example, we describe below the depth estimation process and a triangulation-based calibration method. VGGT is a feed-forward neural network capable of predicting depth maps and tracking 2D keypoints across frames. We input six surrounding camera views into VGGT to generate per-view depth predictions. Following the original VGGT setup, the input images are resized to a resolution of 294 $\times$ 518. Although VGGT produces consistent and high-quality depth estimates across views, the predictions are in an unscaled unit space and do not correspond directly to real-world metric distances. To address this limitation, we leverage VGGT’s built-in 2D point tracking functionality across multiple views at the same time step. Specifically, we select three adjacent cameras including front, front-left, and front-right, and use VGGT to track sparse 2D keypoints across them. By filtering out low-quality matches using the predicted visibility and confidence scores, we obtain reliable point correspondences between camera pairs, as illustrated in Fig. 7. We then triangulate these matched 2D points using the ground-truth camera intrinsics and extrinsics provided by the dataset, resulting in a sparse but metrically accurate 3D point cloud. Finally, we compare the magnitudes of the triangulated 3D points with those reconstructed from the predicted depth maps at the corresponding image locations, and compute a global scaling factor to align the depth predictions with real-world scale. An example of the final scaled depth prediction is shown in Fig. 8.
@@ A.2 Open-Vocabulary Semantic Segmentation
We now describe the process of open-vocabulary semantic segmentation using OpenSeeD [51] as an example. We adopt OpenSeeD primarily to ensure a fair comparison with SelfOcc [12]. Nevertheless, our system is loosely coupled with VFMs and fully compatible with more advanced ones, such as REX-Omni [17]. As shown in Fig. 9, OpenSeeD’s predictions often exhibit noisy and unclear boundaries. Since our focus is not prompt engineering and to ensure a fair comparison with SelfOcc [12], we adopt the same query set including: "barrier", "bicycle", "bus", "car", "construction_vehicle", "crane", "motorcycle", "person", "traffic_cone", "trailer", "trailer_truck", "truck", "road", "sidewalk", "terrain", "grass", "building", "wall", "tree", "sky".
@@ A.3 Tracking with RAFT for TT-OccCamera
For $\operatorname{TT\text{-}OccCamera}$, we estimate the optical flow $F_{\text{opt}}$ between two consecutive frames from the same camera using RAFT [34]. We then compute the ego-motion-induced flow $F_{\text{ego}}$ based on the ground-truth camera intrinsics and extrinsics of the adjacent frames, along with the predicted depth from 3DVFMs. By subtracting the ego flow from the observed optical flow, we obtain the dynamic flow $F_{\text{dyn}}=F_{\text{opt}}-F_{\text{ego}}$, which theoretically captures the motion of dynamic objects in the environment. Although this 2D dynamic flow could, in principle, guide the 3D motion of dynamic Gaussians, back-projecting it into 3D space tends to amplify errors from RAFT and 3DVFMs, resulting in unstable Gaussian motion. To mitigate this, we adopt a compromise strategy by thresholding the dynamic flow magnitude to obtain a dynamic mask that identifies likely moving regions. In the ideal case, a simple thresholding on the magnitude of $F_{\text{dyn}}$ would yield a reliable binary mask for dynamic regions. However, since both $F_{\text{opt}}$ and $F_{\text{ego}}$ are derived from 2D estimations and are subject to noise and inaccuracies, the resulting $F_{\text{dyn}}$ is often highly unreliable and noisy. To further refine the dynamic flow, we leverage the the cues from semantic segmentation models, which provides relatively cleaner object boundaries, to refine the dynamic flow magnitude map. As illustrated in Fig. 10, the raw dynamic flow is noisy, and thresholding it directly often produces fragmented masks that do not correspond to coherent objects. After incorporating instance masks from segmentation models, high-magnitude errors on the background are suppressed, and the resulting dynamic masks become more object-aligned, either an entire object is identified as dynamic or it is not, effectively eliminating partial or spurious activations. The corresponding 3D Gaussians projected onto these regions are treated as dynamic and excluded from static accumulation in the next frame. While this approach does not allow accumulation of dynamic objects as in the LiDAR-based variant, it effectively reduces artifacts caused by noisy motion cues and temporal inconsistencies.
@@ A.4 Tracking with LiDAR for TT-OccLiDAR
Tracking in $\operatorname{TT\text{-}OccLiDAR}$ is generally more reliable than in $\operatorname{TT\text{-}OccCamera}$, as LiDAR point clouds provide more accurate and consistent geometric information. We follow a straightforward strategy: cluster first, then align via ICP. First, we project LiDAR points onto the instance masks predicted by the segmentation model, thereby associating each point with a specific foreground object. Due to the often imprecise boundaries of predicted masks, the resulting instance-level point sets can contain substantial noise. To address this, we apply DBSCAN clustering [9] to each instance’s point cloud to extract its core structure and eliminate outliers. This approach proves effective in significantly removing noise, as illustrated in the left column of Fig. 11, where gray points are obtained by directing projecting onto segmentation masks and green points represent the denoised output after DBSCAN clustering (slightly translated for observation). We then perform object-level matching across adjacent frames based on the spatial proximity and shape similarity of the filtered point clusters. For each matched pair, the 3D flow is estimated using the Iterative Closest Point (ICP) algorithm [1]. Qualitative results are presented in the right column of Fig. 11, where green, blue, and red points represent the source points, destination points, and the ICP-transformed source points, respectively. Green arrows indicate the estimated 3D flow vectors. The effectiveness of the ICP-based alignment can be clearly observed. Finally, matched points are propagated to the next frame, while unmatched instances from the previous frame are discarded to avoid the accumulation of errors caused by moving or disappearing objects.
@@ B.1 Performance under Challenging Conditions
A key strength of our system is that it is built upon large-scale foundation models, which are trained on diverse datasets and therefore exhibit strong generalization beyond specific domains. Leveraging these models gives TT-Occ a natural advantage over task-specific approaches such as SelfOcc [12], particularly under challenging conditions. To assess robustness, we evaluated both the camera- and LiDAR-based variants on nighttime and rainy scenes from the nuScenes [5] dataset. As shown in Table 6, TT-Occ consistently surpasses SelfOcc across all challenging scenarios, demonstrating improved accuracy and robustness.
Table 6: Robustness evaluation on rainy and nighttime nuScenes [5] scenes (mIoU). R: rainy. N: nighttime. TT-Occ variants consistently outperform the task-specific SelfOcc [12] model across all challenging conditions.
Scene ID
@@
$\mathrm{TT\text{-}OccCamera}$ 21.2 13.6 11.2 9.4 10.3 13.1
$\mathrm{TT\text{-}OccLiDAR}$ 27.0 18.9 12.0 13.9 16.2 17.6